# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### v1

@@ IV-F Experiment Setup - VideoCLIP
This framework, referred to as VideoCLIP, is an adaptation of the frame-based CLIP model. It incorporates a 3D Convolutional Neural Network (CNN) backbone, specifically designed for video processing. For additional information, readers can refer to [40]. The key difference in this model is its consideration of the temporal information of video data when making a prediction. When the model needs to make a prediction about a particular frame, it does not just look at that single frame. Instead, it combines 14 frames that come before and 15 frames that come after the target frame to form a small clip. A total of 30 frames are used together for prediction. This video piece is fed into the pre-trained CLIP backbone and finally, majority voting is conducted to produce the final prediction. Fig. 1 shows an overview of the VideoCLIP model pipeline. This technique enables the model to consider not only the content of the targeted frame but also the preceding and subsequent frames. In this way, the network can perform classification with more complete temporal information. In this experiment, the learning rate and batch size are set as 0.0003 and 5, respectively, and Adam [41] optimizer is used. Similar to above, we split the drivers into two non-overlapping sets (training and testing) to evaluate the performance on never-before-seen drivers. Finally, we perform seven-fold cross-validation on the data with sampling rates of 1,2,5 and 20 frames per second.
IV-G Experiment Setup for CLIP vs. Other Models
@@
For performance comparison, we further trained several mainstream CNN backbones, which are commonly used for distracted driving recognition. We considered five different backbones, namely AlexNet [42], VGG16 [16], MobileNetv2 [43], ResNet18 and ResNet50 [25], and used them on frames sampled from videos in SynDD1 dataset. We used ImageNet pre-trained weights and fine-tuned the models with a learning rate of 0.001 and a batch size of 64. We used Adam [41] as the optimizer. We report the experimental results and outcome of our observation in Section V.
@@ V Results
V-A Performance of CLIP-based Frameworks
We present the performance of various CLIP-based frameworks on two distracted driving datasets in Table II. On the SynDD1 dataset, VideoCLIP with an optimal hyperparameter setting demonstrated superior performance with a Top-1 accuracy of 81.85% and a Top-3 accuracy of 91.96% for distracted driving recognition. This was followed by the Multi-frameCLIP framework with an optimal hyperparameter setting. Single-frameCLIP and Zero-shotCLIP reported lower Top-1 accuracies on the SynDD1 dataset compared to the Statefarm dataset.
TABLE II: Performance of CLIP-based
@@ V-A Performance of CLIP-based Frameworks
Framework SynDD1 StateFarm
Top-1 Top-3 Top-1 Top-3
Zero-shotCLIP 54.00 74.50 58.22 67.63
Single-frameCLIP 57.35 82.25 83.15 95.75
Multi-frameCLIP 70.36 75.38 - -
VideoCLIP 81.85 91.96 - -
@@
For the zero-shot experiments, we test the Zero-shotCLIP model on the distracted driving dataset without any prior training. The model encounters previously unseen frames extracted from distracted driving datasets. The RGB input images are resized to a resolution of 224x224x3. During the zero-shot evaluation, the framework extracts visual and textual features from the input images and corresponding textual prompts, respectively. Subsequently, the cosine similarity between these visual and textual features is computed, resulting in a ranked list of the most probable classes based on their similarity scores. The zero-shot transfer result for eight classes on the SynDD1 dataset is presented in Table III for different CLIP backbones. It can be seen that the ViT-L/14@336px vision transformer-based Zero-shotCLIP outperforms other variants in terms of Top-1 accuracy. It should be noted that the ViT-L/14@336px backbone processes input images at a higher RGB resolution (336x336x3). The key observation from these results is that Zero-shotCLIP demonstrates 74.5% Top-3 accuracy in recognizing distracted driving without requiring dataset-specific training.
TABLE III: Zero-shot transfer on SynDD1 dataset
CLIP visual backbone Top-1 accuracy Top-3 accuracy
ViT-L/14 53.50 74.50
ViTL/14@336px 54.00 65.00
ViT-B/16 47.50 64.50
ViT-B/32 38.00 45.50
@@
We present the performance of the Single-frameCLIP model with optimal hyperparameter setting on the StateFarm and SynDD1 datasets in Table IV. Single-frameCLIP results showcase the advantage of task-based fine-tuning. For the StateFarm dataset, the Single-frameCLIP model with the CLIP-L/14 backbone achieved a Top-1 accuracy of 81.85%. On the SynDD1 dataset, the same model reached a Top-1 accuracy of 57.35%. The results further confirm the robustness and effectiveness of the CLIP-L/14 backbone in the Single-frameCLIP model for the task of distracted driving action recognition. It should be noted that the task’s complexity and the specifics of the datasets may account for the difference in accuracy scores between the two datasets.
@@
Single-frameCLIP Dataset Camera-view Top-1
backbone accuracy
CLIP-L/14 StateFarm Dashboard 83.15
CLIP-L/14 SynDD1 Dashboard 57.35
@@
Our Multi-frameCLIP framework involves employing majority voting over a set of frames to determine the final prediction of an action class in a video. We investigate various sampling rates of 0.5, 1, 2, 5, and 20 frames per second, subject-level cross-validation along with different camera views and CLIP backbones. The result with the optimal setting is presented in Table V. The best perfomance on the SynDD1 dataset is achieved with the CLIP-L/14 backbone from the dashboard camera-view footage sampled at 1 frame per second. This setting provided a Top-1 accuracy of 70.36%. We have observed that the model achieves the peak performance at a sampling rate of 1 FPS. Any increase in the sampling rate resulted in a decline in accuracy (as shown in Table X), since the inclusion of similar-looking frames acts as a form of noisy signal in the majority voting pool. On the other hand, reducing the sampling rate below 1 compromised performance, since critical temporal information that is essential for final action prediction gets overlooked. Therefore, we can conclude that 1 FPS is the optimal sampling rate for the Multi-frameCLIP model to adequately capture relevant temporal information. The main advantage of this model is that it processes several frames simultaneously, exploiting the temporal information between consecutive frames. This approach, combined with the high-level semantic understanding of the CLIP model, resulted in an improved performance, as evidenced by the Top-1 accuracy. The optimal hyperparameters were determined through a series of experiments to optimize the model’s overall performance.
@@
Multi-frameCLIP Dataset FPS Camera-view Top-1
backbone accuracy
CLIP-L/14 SynDD1 1 Dashboard 70.36
@@
Similar to Multi-frameCLIP, we conducted a series of tests to obtain the optimal hyperparameter setting for the VideoCLIP model and report the results obtained on the SynDD1 dataset in Table VI. The best performance of 81.85% Top-1 accuracy is achieved with the VideoCLIP built on the CLIP-L/14 backbone and with a sampling rate of 20 frames per second. The performance achieved by the VideoCLIP model validates its efficacy in understanding distracted actions from driving video footage. We have observed that the sampling rate has a significant effect on the final accuracy. Moreover, in contrast to Multi-frameCLIP, VideoCLIP uses a sliding window of 30 frames, and employs a majority voting scheme on the outputs of these small video clips. In our experiments, the VideoCLIP network achieved the best performance when the sampling rate was 20 FPS. If the sampling rate is too small, e.g. 1 FPS, there will be one frame sampled from every second of an activity. Since VideoCLIP uses 30 frames to build a clip, this clip will span 30 seconds which can be longer than the action and confuse the network. For example, an action video lasting 20 seconds should contain 20 frames, however, in these 20 frames, only 10 frames contain the valid information of that action for prediction. If we sampled 30 frames from those 20 frames to make up a window as the input to the network, most frames in the window are noise. In addition, the sliding window approach in VideoCLIP provides a more holistic representation of the video data, which leads to improved prediction accuracy. In the following sections, we will delve into the performance of our models under real-world variabilities and scrutinize the model’s efficacy across diverse scenarios.
@@
VideoCLIP Dataset FPS Camera-view Top-1
backbone accuracy
CLIP-L/14 SynDD1 20 Dashboard 81.85
@@ V-B Performance Comparison with Other Works
In this section, we compare the performance of the Single-frameCLIP model with traditional distracted driving backbones, namely ResNet18, Mobilenetv2, and AlexNet, on both StateFarm and SynDD1 datasets. The Top-1 accuracy values are presented in Tables VIIVIII and Fig. 6. It should be noted that, across all model comparisons in our experiments, we maintain consistency by using the dashboard camera view (which yielded the best results) along with the same action classes. The detailed experiment results for ten distracted driving action classes on the StateFarm dataset are shown in Table VII, where the first column entries represent the drivers used in the testing split. This data split involved leaving three drivers out for testing. For instance, p002p014p012 signifies that drivers with IDs p002, p014, and p012 were set aside for the first fold. The remaining drivers were used for the training set. We follow this procedure eight times, each time keeping different groups of drivers separate for testing. We conduct 8-fold driver-level cross-validation on the StateFarm dataset. The other columns display the Top-1 accuracy of each model for different test splits. The last row provides an average Top-1 accuracy for each model across all folds. Amongst all the models, the Single-frameCLIP model with ViT-L/14 architecture emerged as the best performing, achieving an average Top-1 accuracy of 83.15%. Fig. 6 visually represents these results. The bars illustrate the average Top-1 accuracy of the models, while the vertical lines on the top indicate the variance in Top-1 accuracies across all eight folds.
@@
Test Drivers ResNet18 Mobilenetv2 AlexNet CLIP-L/14
@@
The comparison results obtained with the eight distracted driving classes in the SynDD1 dataset are presented in Table VIII. Similar to the experiments on the StateFarm dataset, this experiment involved 7-fold driver-level cross-validation leaving two drivers out for testing in each fold. The third column in Table VIII represents the average Top-1 accuracy across all the 7-folds. The results reveal that the CLIP-L/14 model outperformed others, achieving a Top-1 accuracy of 57.347% (average accuracy across 7-folds). The remaining models followed with lower accuracies, with CLIP-B/16 achieving 48.940%, and CLIP-RN101 at 48.521%. In conclusion, the Single-frameCLIP model with the ViT-L/14 architecture demonstrated superior performance in terms of Top-1 accuracy both in the StateFarm and SynDD1 datasets. In contrast, traditional models fell short in achieving performance close to that of the CLIP-L/14 model.
TABLE VIII: Single-frameCLIP-backbones vs. traditional models:Top-1 accuracy on SynDD1 dataset
Average
Models Backbones Top-1
accuracy
Single-frameCLIP CLIP-L/14 57.347
CLIP-B/16 48.940
CLIP-RN101 48.521
CLIP-B/32 35.382
@@
In an effort to analyze the performance of the Single-frame CLIP model with smaller amounts of training data, a series of experiments were conducted utilizing a progressively reduced training dataset. In this experiment, we train the traditional models and our proposed CLIP-based frameworks in a similar experimental setting. From the StateFarm dataset, we separate 3 drivers (drivers with ID p012, p014, p021) for testing and progressively removed 4 drivers from the training set to create a smaller training set, which is $80\%$, $60\%$, $40\%$, and $20\%$ of the original training set. For this process, instead of randomly removing frames, we reduce the training data based on the driver ID, since in real-world scenarios the model will be expected to generalize its learning to new, unseen drivers. The percentage of training data, the number of drivers present in those splits, and the number of samples are shown in Table  IX. For this experiment on the StateFarm dataset, we employ an 8-fold cross-validation scheme where we excluded three drivers (driver ID p012, p014, p021) from every fold and used them as test subjects. We picked CLIP-ViT/L-14 backbone for the Single-frameCLIP. All the models used the same training samples and were finally tested on the same test set. The performances are presented in Fig. 7. The results suggest that reducing the size of the training set does not necessarily compromise the Single-frameCLIP model’s performance. Unlike the other models assessed, which displayed a decrease in accuracy, the Single-frameCLIP model maintained its performance levels consistently even when the quantity of training data was reduced to 20% of the original amount.
@@
Figure 7: Performance comparison of the Single-frameCLIP model and traditional models trained on varying proportions of the StateFarm dataset. Training sets of 80%, 60%, 40%, and 20% were created by progressively removing drivers based on their ID. The figure reports the models’ average accuracies obtained through an 8-fold cross-validation scheme based on driver ID.
@@
We present the Top-1 accuracy results of Multi-frameCLIP and traditional models on the SynDD1 dataset in Table X. The evaluated models include several versions of CLIP backbones, ResNet18, ResNet50, Modilenetv2, AlexNet, and Vgg16. The results are presented for different sampling rates and dashboard camera view as this view yielded the best results. From the table, we can see that the highest Top-1 accuracy of 70.36% is achieved by CLIP-L/14 with a sampling rate of 1 FPS. On the lower end, AlexNet and Vgg16 appear to perform the worst across all sampling rates, with accuracy rates hovering around 12.62%. In general, the accuracy appears to vary depending on the model and the sampling rate (discussed in the following sections), indicating that these are significant factors in performance on the SynDD1 dataset.
TABLE X: Multi-frameCLIP at different sampling rates on SynDD1 dataset
Sampling Rate CLIP-L/14 CLIP-B/16 CLIP-RN101 CLIP-B/32 ResNet18 ResNet50 Modilenetv2 AlexNet Vgg16
@@
TABLE XI: VideoCLIP at different sampling rates on SynDD1 dataset

### v3

@@ IV-H Experiment Setup - VideoCLIP
The VideoCLIP is an adaptation of the frame-based CLIP model. It incorporates a 3D Convolutional Neural Network (CNN) backbone specifically designed for video processing. For additional information, readers can refer to [67]. The key difference in this model is its consideration of the temporal information of video data when making a prediction. When the model needs to make a prediction about a particular frame, it does not just look at that single frame. Instead, it combines 14 frames that come before and 15 frames that come after a particular target frame to form a small video footage. A total of 30 frames (including the target frame) are used together for prediction. This video piece is fed into the pre-trained CLIP backbone and a majority voting step is conducted to produce the final prediction. Fig. 1 shows an overview of the VideoCLIP model pipeline. This technique enables the model to consider not only the content of the targeted frame but also the preceding and subsequent frames. In this way, the network can perform classification with more complete temporal information. In this experiment, the learning rate and batch size are set as 0.0003 and 5, respectively, and Adam [68] optimizer is used. We followed the ‘leave a set of driver out’ method as before and split the drivers into two non-overlapping sets (training and testing) to evaluate the performance of never-before-seen drivers. Finally, we perform a 7-fold cross-validation on the splits with sampling rates of 1,2,5 and 20 frames per second.
IV-I Experiment Setup for CLIP vs. Other Models
@@
For performance comparison, we further trained several mainstream CNN backbones, which are commonly used for distracted driving recognition. For this experiment, we considered five different backbones - AlexNet [43], VGG16 [41], MobileNet-v2 [45], ResNet18 and ResNet50 [42], and implemented them on the frames sampled from the video data. We used the ImageNet pre-trained weights and fine-tuned the models with a learning rate of 0.001 and a batch size of 64. We used Adam [68] as the optimizer. We report the experimental results and outcome of our observation in Section V.
@@ V Results
V-A Performance of the Proposed CLIP-based Frameworks
We present the performance summary of various CLIP-based frameworks on the four distracted driving datasets in Table II. The VideoCLIP model with an optimal hyperparameter setting demonstrated the most superior performance with a Top-1 accuracy of 81.85%, 98.44% and 97.86% on SynDD1, DMD and SAM-DD datasets, respectively. This was followed by the Multi-frameCLIP framework with an optimal hyperparameter setting. Single-frameCLIP and Zero-shotCLIP reported lower Top-1 accuracies on the SynDD1 dataset compared to the Statefarm dataset.
TABLE II: Performance Summary of CLIP-based
@@ V-A Performance of the Proposed CLIP-based Frameworks
Models SynDD1 StateFarm DMD SAM-DD
(ours) Top-1 Top-3 Top-1 Top-3 Top-1 Top-3 Top-1 Top-3
Zero-shotCLIP 54.00 74.50 58.22 67.63 48.12 79.56 66.52 86.33
Single-frameCLIP 57.35 82.25 83.15 95.75 82.65 97.76 89.14 98.78
Multi-frameCLIP 70.36 75.38 N/A N/A 93.39 95.74 89.14 90.27
VideoCLIP 81.85 91.96 N/A N/A 98.44 99.80 97.86 99.76
@@
For the zero-shot experiments, we test the Zero-shotCLIP model on the distracted driving dataset without any prior training. The model encounters previously unseen frames extracted from distracted driving datasets. The RGB input images are resized to a resolution of 224x224x3. During the zero-shot evaluation, the framework extracts visual and textual features from the input images and corresponding textual prompts. Subsequently, the cosine similarity between these visual and textual features is computed, resulting in a ranked list of the most probable classes based on their similarity scores. The zero-shot transfer results are presented in Table III for different CLIP backbones. It can be seen that the ViT-L/14@336px vision transformer-based Zero-shotCLIP outperforms other variants in terms of Top-1 accuracy. It should be noted that the ViT-L/14@336px backbone processes input images at a higher RGB resolution (336x336x3). A key observation from this table is that ViT-L/14@336px based Zero-shotCLIP reached up to 86.33% Top-3 accuracy in the SAM-DD dataset without requiring dataset-specific training.
TABLE III: Zero-shot transfer results
CLIP SynDD1 DMD SAM-DD
backbones Top-1 Top-3 Top-1 Top-3 Top-1 Top-3
ViT-L/14 53.50 74.50 47.97 79.56 62.51 84.52
ViT-L/14@336px 54.00 65.00 48.12 77.16 66.52 86.33
ViT-B/16 47.50 64.50 31.29 54.95 18.67 79.18
ViT-B/32 38.00 45.50 28.39 47.76 10.48 40.09
@@
We demonstrate the performance of the Single-frameCLIP model with optimal hyperparameter setting on the four datasets in Table IV. Single-frameCLIP results showcase the advantage of task-based fine-tuning over Zero-shotCLIP. For the SAM-DD dataset, the Single-frameCLIP model with the ViT-L/14 backbone achieved a Top-1 accuracy of 89.14% with the side-view frames. On the SynDD1 dataset, the same model reached a Top-1 accuracy of 57.35%. The results further confirm the robustness and effectiveness of the ViT-L/14-based backbones in the Single-frameCLIP model for distracted driving recognition tasks. It should be noted that the ‘leaving a set of driver out’ and the $k$-fold cross-validation method played a vital role in determining the optimal hyperparameter setting.
@@
Single-frameCLIP Dataset Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 StateFarm Side-view 83.15
ViT-L/14 SynDD1 Dashboard 57.35
ViT-L/14@336px DMD Side-view 82.65
ViT-L/14 SAM-DD Side-view 89.14
@@
The Multi-frameCLIP framework involves employing majority voting over a set of consecutive frames to determine the final prediction of an action in a video clip. We investigate various sampling rates of 0.5, 1, 2, 5, and 20 frames per second, subject-level cross-validation along with different camera views and CLIP backbones. The result with the optimal setting is presented in Table V. The best performance is achieved with the ViT-L/14@336px backbone on the DMD dataset for the side-view video clip sampled at 1 frame per second. This optimal setting provided a Top-1 accuracy of 93.39%. We have observed that the model achieves peak performance at a sampling rate of 1 FPS. Any increase in the sampling rate resulted in a decline in accuracy (as shown in Table XIII) since the inclusion of similar-looking frames acts as a form of noisy signal in the majority voting pool. On the other hand, reducing the sampling rate below 1 compromised the performance since critical temporal information that is essential for final action prediction gets overlooked. Therefore, we can conclude that 1 FPS is the optimal sampling rate for the Multi-frameCLIP model to adequately capture relevant temporal information. The main advantage of this model is that it processes several frames simultaneously, exploiting the temporal information between two consecutive frames. This approach, combined with the high-level semantic understanding of the CLIP model, resulted in an improved performance. The optimal hyperparameters were determined through a series of experiments to optimize the model’s overall performance.
@@
Multi-frameCLIP Dataset FPS Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 SynDD1 1 Dashboard 70.36
ViT-L/14@336px DMD 1 Side-view 93.39
ViT-L/14 SAM-DD 1 Side-view 89.14
@@
VideoCLIP Dataset FPS Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 SynDD1 20 Dashboard 81.85
ViT-L/14@336px DMD 1 Side-view 98.44
ViT-L/14 SAM-DD 1 Side-view 97.86
Similar to Multi-frameCLIP, we conduct a series of tests to obtain the optimal hyperparameter setting for the VideoCLIP model and report the results in Table VI. The best performance of 97.86% Top-1 accuracy is achieved on the SAM-DD dataset with the ViT-L/14 backbone and with a sampling rate of 1 frame per second. The performance achieved by the VideoCLIP model validates its efficacy in understanding distracted actions from driving video footage. We have observed that the sampling rate has a significant effect on the final accuracy. Moreover, in contrast to Multi-frameCLIP, VideoCLIP uses a sliding window of 30 frames and employs a majority voting scheme on the outputs of these small video clips. It is important to note that on the SynDD1 dataset, the VideoCLIP achieved the best performance at 20 FPS sampling rate. If the sampling rate is too small, e.g., 1 FPS, there will be one frame sampled from every second of an activity. Since VideoCLIP uses 30 frames to build a video clip, this clip will span 30 seconds which can be longer than the average action duration (from Fig. 5 the average action duration for SynDD1 dataset is $\sim$18.7 sec) and confuse the network. For example, an action video lasting 20 seconds should contain 20 frames at 1FPS; however, in these 20 frames, only 10 frames contain valid information about that action. Now, if we sample 30 frames from those 20 frames to make up a window as the input to the VideoCLIP network, most frames in that window are noise. In addition, the sliding window approach in VideoCLIP provides a more holistic representation of the video data, which leads to improved prediction accuracy. In the following sections, we will delve into the performance of our models under real-world variabilities and scrutinize the model’s efficacy across diverse scenarios.
@@ V-B Performance Comparison with Other Works
In this section, we compare the performance of the Single-frameCLIP model with traditional distracted driving backbones on frame-based datasets. We provided two distinct tables to highlight how the accuracy scores change across different test driver sets and model backbones on the StateFarm and SynDD1 datasets. The Top-1 accuracy scores from our runs are presented in Tables VIIVIII and Fig. 6. It should be noted that, across all model comparisons in our experiments, we maintain consistency by using the same experimental settings along with the same action classes. Table VII shows the detailed breakdown of our ‘leave a set of drivers out’ approach that we used for testing on the StateFarm dataset. The first column entries represent the drivers included in the testing split. This data split involved leaving three drivers out for testing. For instance, ‘p002p014p012’ signifies that drivers with driverIDs ‘p002’, ‘p014’, and ‘p012’ were considered for the test set in the first fold. The remaining drivers were used for the training set. We follow this procedure eight times, each time keeping different groups of drivers separate for testing. We conduct 8-fold driver-level cross-validation on the StateFarm dataset. The other columns display the Top-1 accuracy of each model for different test splits. The last row provides an average Top-1 accuracy for each model across all folds. Amongst all the models, the Single-frameCLIP model with ViT-L/14 architecture emerged as the best performing, achieving an average Top-1 accuracy of 83.15%. Fig. 6 visually represents these results. The bars illustrate the average Top-1 accuracy of the models, while the vertical lines on the top indicate the variance in Top-1 accuracies across all eight folds.
@@
Test DriverID ResNet18 MobileNet-v2 AlexNet ViT-L/14
@@
TABLE VIII: Single-frameCLIP vs. traditional modelsTop-1 accuracy on SynDD1 dataset
Average Leaving
Models Backbones Top-1 a set of
accuracy driver out
Single-frameCLIP ViT-L/14 57.347
ViT-B/16 48.940
RN101 48.521
ViT-B/32 35.382
@@
The results obtained with the eight distracted driving classes in the SynDD1 dataset are presented in Table VIII. Similar to the experiments on the StateFarm dataset, this experiment involved 7-fold driver-level cross-validation with ‘leaving two drivers out’ for testing in each fold. The third column in Table VIII represents the average Top-1 accuracy across all the 7-folds. The results reveal that the ViT-L/14 model outperformed others, achieving a Top-1 accuracy of 57.347% (average accuracy across 7-folds). The remaining models followed with lower accuracies, with ViT-B/16 achieving 48.940% and RN101 at 48.521%. In conclusion, the Single-frameCLIP model with the ViT-L/14 architecture demonstrated superior performance in terms of Top-1 accuracy both in the StateFarm and SynDD1 datasets.
@@
In an effort to analyze the performance of the Single-frame CLIP model with smaller amounts of training data, a series of experiments were conducted utilizing a progressively reduced training dataset. In this experiment, we train the traditional models and our proposed CLIP-based frameworks in a similar experimental setting on the StateFarm dataset. We separate 3 drivers (driverID p012, p014, p021) for testing and progressively removed four drivers from the training set to create a smaller training set, which is $80\%$, $60\%$, $40\%$, and $20\%$ of the original training set. For this process, instead of randomly removing frames, we reduce the training data based on the driverID since, in real-world scenarios, the model will be expected to generalize its learning to new, unseen drivers. The percentage of training data, the number of drivers present in those splits, and the number of samples are shown in Table  IX. For this experiment on the StateFarm dataset, we employ an 8-fold cross-validation scheme where we excluded three drivers (driverID p012, p014, p021) from every fold and used them as test subjects. We picked the ViT/L-14 backbone for the Single-frameCLIP. All the models used the same training samples and were finally tested on the same test set. The outcomes are presented in Fig. 7. The results suggest that reducing the size of the training set does not necessarily compromise the Single-frameCLIP model’s performance. Unlike the other models assessed, which displayed a decrease in accuracy, the Single-frameCLIP model maintained its performance levels consistently even when the quantity of training data was reduced to 20% of the original amount.
@@
Figure 7: Performance comparison of the Single-frameCLIP model and traditional models trained on varying proportions of the StateFarm dataset. Training sets of 80%, 60%, 40%, and 20% were created by progressively removing drivers based on their driverID. The figure reports the models’ average accuracies obtained through an 8-fold cross-validation scheme based on driverID.
@@
TABLE X: Multi-frameCLIP vs. traditional models(at Sampling rate 1FPS)
Video Models Top-1 Modality Leaving a set
Datasets Acc.(%) of driver out
DMD AlexNet 45.86 RGB
VGG16 56.15 RGB
MobileNet-v2 85.77 RGB
ResNet18 85.95 RGB
Multi-frameCLIP 93.39 RGB
SAM-DD VGG16 86.93 RGB
AlexNet 88.50 RGB
MobileNet-v2 92.95 RGB
Multi-frameCLIP 89.14 RGB
SynDD1 AlexNet 12.56 RGB
VGG16 12.56 RGB
MobileNet-v2 33.65 RGB
ResNet50 43.49 RGB
ResNet18 47.55 RGB
Multi-frameCLIP 70.36 RGB
We provide an extensive performance comparison of Multi-frameCLIP with traditional models on three different datasets in Table XIII. To ensure consistency and reliability in our experiments, we run the listed models at a 1FPS sampling rate with the same 7-fold cross-validation and ‘leaving a set of driver out’ settings across all the datasets. The Multi-frameCLIP model shows superior performance compared to the other baselines, achieving a Top-1 accuracy of 93.39% on DMD, 89.14% on SAM-DD, and 70.36% on the SynDD1 dataset. These results are significantly higher than those of models such as AlexNet, VGG16, MobileNet-v2, and ResNet variants when tested under the same conditions. In addition, it showcases the Multi-frameCLIP model’s efficacy in capturing driver behavior in different scenes, making it a suitable benchmark for distracted driving detection applications. However, it is important to note that accuracy fluctuates with changes in hyperparameters such as model type, sampling rate, camera view, and so on (discussed in the following sections), indicating that these are significant factors to consider to get an optimized performance.
4)  VideoCLIP vs. Traditional Models
We showcase a comparative analysis of the VideoCLIP against the traditional baselines in Table XI. It shows that VideoCLIP gets significant performance gain and outperforms its counterparts across three different datasets. On the DMD dataset, VideoCLIP outperforms all other models, including advanced configurations like MobileNet-v2 (MLP), 3D CNN variants and modalities, achieving a Top-1 accuracy of 98.44%. Similarly, in the SAM-DD dataset, VideoCLIP tops the chart with 97.86% accuracy, surpassing other SOTA models such as Swin Transformers and ViTs. Furthermore, on the SynDD1 dataset, VideoCLIP maintains its lead with an 81.85% accuracy at a 20FPS rate, which is substantially higher than the average Top-1 accuracy of the competing model [69]. These results highlight the efficacy of the VideoCLIP model in harnessing RGB modality and underscore its robustness across different datasets within the ’leave a set of driver out’ setting.
TABLE XI: VideoCLIP vs. traditional models(with Sampling rate 1FPS)
Video Models Top-1 Modality Leaving a set
Datasets Acc.(%) of driver out
DMD ShuffleNet-v2 84.80 RGB x
3D-ShuffleNet-v2 88.50 RGB x
MobileNet-v2 89.90 RGB x
MobileNet-v2 85.77 RGB
3D-MobileNet-v2 90.30 RGB x
MobileNet-v2(MLP) 91.60 RGB x
Inception-v3 91.80 RGB x
MobileNet-v2(MLP) 93.70 RGB+IR+D x
MobileNet-v2(MLP) [35] 93.90* IR+D x
VideoCLIP 98.44 RGB
SAM-DD MobileNet-v3 94.78 RGB
ShuffleNet-v2 95.56 RGB
ResNet18 95.91 RGB
VGG16 96.86 RGB
ViT 97.09 RGB
Swin-T [4] 97.33* RGB
VideoCLIP 97.86 RGB
SynDD1 VideoMAE [69] 71.35** RGB
VideoCLIP(1FPS) 74.64 RGB
VideoCLIP(20FPS) 81.85 RGB
*best performing model proposed by the author
**average Top-1 accuracy for the same classes used in VideoCLIP
TABLE XII: Single-frameCLIP: F1-score on DMD and SAM-DD
Action F1-score F1-score
class DMD SAM-DD
drinking water 0.96 0.98
hair and makeup 0.91 0.84
talking to phone 0.97 0.96
adjusting radio 0.86 -
reaching to backseat 0.68 0.91
reaching to side 0.58 -
driving safely 0.68 0.93
talking to passenger 0.87 -
texting on phone 0.92 0.84
yawning 0.76 -
TABLE XIII: Multi-frameCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)
Sampling Rate ViT-L/14 ViT-B/16 RN101 ViT-B/32 ResNet18 ResNet50 Modilenetv2 AlexNet VGG16
@@
TABLE XIV: VideoCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)